# Outlier Gradient Analysis: Efficiently Improving Deep Learning Model Performance via Hessian-Free Influence Functions

### v2

@@ 8 Conclusion
In this paper, we addressed two key challenges associated with the application of influence estimation functions in deep models—the assumptions of model/loss convexity and the computational demands for inverting the Hessian matrix. Focusing on the data-centric task of trimming detrimental samples, we established an equivalence transformation between influence functions and outlier analysis in the gradient space. Our approach relaxes the convexity assumption and extends influence estimation to non-convex deep models. Through comprehensive experiments on synthetic datasets and various application domains (code details in Appendix D), including noisy label correction for vision models, data selection for NLP models, and influential data identification in LLMs, we demonstrated that outlier gradient analysis and trimming outperformed many existing influence-based approaches in deep learning scenarios, highlighting its advantages as a method for influence estimation.
@@
Appendix A Detailed Information on Datasets and Model Training
We describe dataset details as well as model training and other information used in the main paper, below.
A.1 Datasets
@@ A.1.1 Synthetic Datasets
We conduct experiments for our proposed outlier gradient analysis and other baselines on two synthetic datasets. The first dataset is linearly seperable for logistic regression classification and consists of 150 training samples and 100 test samples. These are created using the scikit-learn library’s make_blobs function. [^3] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_blobs.html. For each of the two binary classes, we manually flip the labels of 10 samples (5 for each class) to add noise to the dataset. The second dataset is the non-linear half moons dataset so that we can train an MLP network with two hidden layers with ReLU activations. The training set has 250 samples and the test set has 100 samples, and the dataset is generated using the scikit-learn library’s make_moons function. [^4] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_moons.html. Here too, we manually flip the labels of 20 samples (10 from each class) to add noise to the data.
A.1.2 CIFAR-10N and CIFAR-100N
@@
Both the CIFAR-10N and CIFAR-100N datasets [Wei et al., 2022] consist of the same input images that make up the CIFAR-10 (10 classes) and CIFAR-100 (100 classes) datasets [Krizhevsky et al., 2009], respectively. Each input is a 32x32 RGB image with dimension (3,32,32). However, for CIFAR-10N and CIFAR-100N, the labels are noisy, as they contain real-world human annotation errors collected using 3 annotators on Amazon Mechanical Turk. As these datasets are based on human-annotated noise, they model noisy real-world datasets more realistically, compared to synthetic data alternatives. The training set for both datasets contains 50000 image-label pairs, and the test set contains 10000 image-label pairs that are free from noise. For CIFAR-10N we utilize three noise settings for experiments in the paper– (1) Worst, which is the dataset version with the highest noise rate (40.21%) as the worst possible annotation label for the image is chosen, (2) Aggregate, which is the least noisy dataset (9.03%) as labels are chosen via majority voting amongst the annotations, and (3) Random which has intermediate noise (17.23%) and consists of picking one of the annotators’ labels. We use the first annotator for the random labels. For CIFAR-100N there is only a single noisy setting (Noisy100) due to the large number of labeling classes, and the overall noise rate is 40.20%.
A.1.3 GLUE Datasets
@@
The GLUE or the General Language Understanding Evaluation [Wang et al., 2018] benchmark datasets consist of a number of benchmarks for training, evaluating, and analyzing natural language models. As in the DataInf paper [Kwon et al., 2024], we utilize the four binary classification subset datasets: QNLI, SST2, QQP, and MRPC for experiments. Here, these datasets cover a wide variety of natural language task domains. For instance QNLI [Wang et al., 2018] covers natural language inference, SST2 [Socher et al., 2013] covers sentiment analysis, QQP  [^5] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs covers question answering, and MRPC [Dolan and Brockett, 2005] covers paraphrase detection. We use the same datasets as in Kwon et al. [2024], where the training and test splits are obtaned from the Huggingface datasets [^6] https://huggingface.co/docs/datasets library. For QQP and SST2 in Kwon et al. [2024] 4500 training samples and 500 test samples were randomly sampled from the full sets, so we utilize these in our experiments for fair comparison.
A.1.4 Sentence Transformations
@@
For this benchmark dataset proposed in [Kwon et al., 2024], the LLM is required to perform a specific transformation on an input sentence. There are 10 different sentence transformations. To help the model learn different transformations, “chatbot” name identifiers are used and each is uniquely associated with each transformation. These are the categories of sentence transformations (taking an example input sentence as “Welcome to the real world.”):
@@
A.1.5 Math With/Without Reasoning
@@ A.2 Models and Methods
We now describe the models and the methods used in our experiments throughout the main paper. First, we describe the ResNet-34 [He et al., 2016] architecture used as the base model for the noisy vision datasets, then the RoBERTa [Liu et al., 2019] NLP transformer model, and then the Llama-2 LLM. [^7] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf. We also describe implementation details and parameter values for the label correction baselines used in Sections 4 and 5 and the influence-based baselines used throughout the paper. Finally, we also describe some key implementation details regarding our outlier gradient analysis approach.
A.2.1 ResNet-34
@@
The ResNet-34 model was proposed in [He et al., 2016] and is a 34 layer convolutional neural network pretrained on the ImageNet-1K dataset at resolution 224 $\times$ 224. The pretrained model block is fine-tuned on the CIFAR-10N/CIFAR-100N training set experiments with default parameters– minibatch size (128), optimizer (SGD), initial learning rate (0.1), momentum (0.9), weight decay (0.0005), and number of epochs (100), for all experiments. Moreover, we directly used the implementation provided by Wei et al. [2022] and made modifications to their code.
A.2.2 RoBERTa
@@
As in [Kwon et al., 2024], we utilize LoRA fine-tuning to fine-tune the RoBERTa-large model, a 355M parameter transformer language model that improves upon the original BERT model in key ways such as implementation and hyperparameter selection. LoRA is applied to every value matrix of the attention layers of the RoBERTa model. The pre-trained model from Huggingface is used. [^8] https://huggingface.co/docs/transformers/model_doc/roberta. A learning rate of 0.0003 and a batch size of 32 is used. The model is fine-tuned over 10 epochs using LoRA and dropout is set to be 0.05 while the rank of the LoRA matrix is set to 4, as recommended in Kwon et al. [2024]. The loss function used is a negative log-likelihood as the datasets are all for binary classification. The LoRA training is enabled using the Huggingface PEFT library. [^9] https://huggingface.co/docs/peft/index. For the influence experiments we have utilized the code provided in [Kwon et al., 2024] and adapted it for our experiments. Moreover, we only compute influences using the training set gradients, and keep the test set hidden from the learning model for fair evaluation.
A.2.3 Llama2-13B-chat LLM
@@
We fine-tune the Llama2 13B parameter instruction tuned LLM using LoRA fine-tuning (applied to every query and value matrix of the attention layer) as in Kwon et al. [2024]. The LoRA parameters are as follows: learning rate is set to be 0.0003, rank of LoRA matrix is set to 8, $\alpha=32$ in 8-bit quantization, and the batch size is set to 32 across 25 fine-tuning epochs. A negative log-likelihood of the generated response is used as the loss function for fine-tuning as before. Here too, we adapt the code provided by Kwon et al. [2024] for our use-cases.
A.2.4 Label Correction Baselines
@@
For each of the label correction baselines employed in Sections 4 and 5Normalized Margin [Northcutt et al., 2021], Self-Confidence [Müller and Markert, 2019], and Confidence-Weighted Entropy [Kuan and Mueller, 2022], we utilize the implementation provided in the Cleanlab [^10] https://github.com/cleanlab/cleanlab/. library. We use default parameters for all three baselines. Note that the baselines are model agnostic and only require predicted labels and associated probabilities for predictions, which we can easily obtain from classifiers.
A.2.5 Influence-Based Baselines
@@
A.2.6 Outlier Gradient Analysis
We now discuss implementation details regarding outlier gradient analysis. Owing to the simplicity of our approach, the implementation is straightforward and follows directly from the algorithm. In most cases, we directly utilize the gradients obtained from the last layer of the model being considered. However, in some cases, the gradient space of samples can be high dimensional. For instance, for CIFAR-100N, the gradient space is of dimension 50000 $\times$ 51200 which unnecessarily increases memory and time complexity of outlier detection. As a result, we reduce the gradient space dimensionality by employing a sparse random projection step [Li et al., 2006] where the reduced dimension is ascertained using the scikit-learn library. [^11] https://scikit-learn.org/. We also utilize sparse random projection in this manner for the Llama-2-13B-chat LLM model experiments to reduce the dimensionality of the gradient space obtained.
Appendix B Additional Results and Experiments
We now provide details on additional experiments. We first provide results for the noisy label datasets and vision models shown in the main paper, but with standard deviation included. Then we conduct experiments showcasing how to select a suitable outlier detection threshold for the iForest algorithm. We also provide experiments on running time of our proposed approach (as well as details on computational complexity), ablation experiments on varying iForest parameters, results on ImageNet, and experiments with ResNet-18 being used as the base model instead of ResNet-34.
@@ Appendix B Additional Results and Experiments
Normalized Margin Northcutt et al. [2021] 91.33 ± 0.11 90.06 ± 0.14 83.57 ± 0.32 60.94 ± 0.59
Self-Confidence Müller and Markert [2019] 91.38 ± 0.19 90.09 ± 0.17 83.65 ± 0.21 60.51 ± 0.51
Confidence Entropy Kuan and Mueller [2022] 91.11 ± 0.34 90.05 ± 0.26 83.63 ± 0.41 60.62 ± 0.26
Gradient Tracing Pruthi et al. [2020] 91.47 ± 0.21 89.98 ± 0.20 83.38 ± 0.58 60.73 ± 0.38
LiSSA Koh and Liang [2017] 91.49 ± 0.34 90.05 ± 0.31 83.38 ± 0.58 60.48 ± 0.29
DataInf Kwon et al. [2024] 91.46 ± 0.17 90.05 ± 0.38 83.40 ± 0.56 60.70 ± 0.31
Self-LiSSA Bejan et al. [2023] 92.07 ± 0.15 89.58 ± 0.11 83.01 ± 0.34 59.48 ± 0.43
@@
B.1 Full Results with Standard Deviation for Vision Model Experiments
In the main paper results of Section 5 we provide accuracy values without the standard deviation listed, due to space constraints. Here, we augment those results by also providing the standard deviation obtained over the 5 runs. These result are denoted in Table 4. It can be seen that the standard deviations are in general low, and overall, outlier gradient trimming has low variance.
B.2 Selecting a Suitable Outlier Removal Threshold $k$
We now discuss why the outlier detection threshold of 5% is suitable for experiments and how it was chosen. Basically, we utilize different values of the trimming budget $k$ ranging from 0% to 12.5% of the training dataset size (in 2.5% intervals), and then measure performance on the unseen test set for the CIFAR-10N dataset. These results are shown in Figure 5. As can be observed, the optimal performance is obtained for the 5% value of $k$, although it is important to note that any trimming undertaken for a $k$ in this range leads to an increase in performance from the original model performance. Owing to these results, we utilize the threshold of 5% for all experiments.
Figure 5: Test set performance measured for different outlier trimming budget values ($k$) as a percentage of the training data.
B.3 Experiments on Running Time and Computational Complexity
We now present running time experiments for outlier gradient analysis on both the CIFAR-10N and CIFAR-100N datasets compared to the other baselines compared in the paper in Table 5. It can be seen that outlier gradient analysis is computationally efficient and a fraction of the original running time of the model. Moreover, it is order of magnitudes faster than the other baselines. Thus, our outlier gradient analysis approach is computationally efficient as an option for trimming detrimental samples and improving model performance. We also provide analytical time complexity comparisons in Table 6. Although, it is important to note that in practice, outlier gradient analysis is much faster than the worst case time complexity, as can be seen in Table 5.
Table 5: Running time results for our outlier gradient analysis approaches and other baselines.

### v3

@@ 9 Conclusion
In this paper, we focused on the key data-centric learning task of identifying detrimental training samples. Influence functions are a leading approach often used for this problem, but possess certain deficiencies when applied to deep models. As part of our proposed solution, we address the key challenge associated with influence functions usage in deep models— the computational demands for inverting the Hessian matrix. This results in a computationally efficient method that possesses high detection accuracy. Our approach, outlier gradient analysis, is based on a conceptual transformation between the influence function formulation and outlier analysis in the gradient space. Through comprehensive experiments on synthetic datasets and various application domains (code details in Appendix E), including noisy label correction for vision models, data selection for NLP models, and even influential data identification in LLMs, we demonstrated that our method outperformed many existing influence-based approaches and baselines in deep learning scenarios.
@@
Appendix A Additional Related Work on Miscellaneous Data-Centric Learning
Many works in the data-centric learning domain study other relevant research questions beyond detrimental sample identification and influence estimation. For instance, datamodels [Ilyas et al., 2022] also estimate training sample contributions, but only for one test sample at a time. Data efficiency approaches [Jain et al., 2023, Paul et al., 2021, Killamsetty et al., 2021] aim to accelerate deep learning training time via subset selection. Data pruning approaches based on novel approximations for leave-one-out influence estimation [Tan et al., 2024] and the model’s generalization gap [Yang et al., 2022] have also been proposed. Model pruning via generalized influence functions has also been studied in Lyu et al. [2023]. Note that after identifying detrimental training samples, one can adopt multiple strategies for recourse. While we focus on removal in this paper, other alternatives could also be used, such as relabeling [Richardson et al., 2023, Kong et al., 2021]. Antidote data augmentation [Chhabra et al., 2022, Li et al., 2023] methods aim to generate synthetic data samples to improve model performance, whereas feature selection approaches [Hall, 1999, Cai et al., 2018] seek to optimize the feature space to only those important for model performance. Active learning [Cohn et al., 1996] methods aim to iteratively identify optimal samples to annotate given a large unlabeled training data pool [Liu et al., 2021, Nguyen et al., 2022, Wei et al., 2015]. Finally, works on poisoning attacks seek to analyze model robustness by perturbing training set samples [Solans et al., 2021, Mehrabi et al., 2021, Chhabra et al., 2023] under natural input constraints. The study of training sample influence has also been extended to recent generative models, such as diffusion models [Dai and Gifford, 2023], through the use of ensembles.
Appendix B Detailed Information on Datasets and Model Training
We describe dataset details as well as model training and other information used in the main paper.
B.1 Datasets
@@ B.1.1 Synthetic Datasets
We conduct experiments for our proposed outlier gradient analysis and other baselines on two synthetic datasets. The first dataset is linearly separable for logistic regression classification and consists of 150 training samples and 100 test samples. These are created using the scikit-learn [Pedregosa et al., 2011] library’s make_blobs function. For each of the two binary classes, we manually flip the labels of 10 samples (5 for each class) to add noise to the dataset. The second dataset is the non-linear half moons dataset so that we can train an MLP network with two hidden layers with ReLU activations. The training set has 250 samples and the test set has 100 samples, and the dataset is generated using the scikit-learn library’s make_moons function. Here too, we manually flip the labels of 20 samples (10 from each class) to add noise to the data.
B.1.2 CIFAR-10N and CIFAR-100N
@@
Both the CIFAR-10N and CIFAR-100N datasets [Wei et al., 2022] consist of the same input images that make up the CIFAR-10 (10 classes) and CIFAR-100 (100 classes) datasets [Krizhevsky et al., 2009], respectively. Each input is a 32x32 RGB image with dimension (3,32,32). However, for CIFAR-10N and CIFAR-100N, the labels are noisy, as they contain real-world human annotation errors collected using 3 annotators on Amazon Mechanical Turk. As these datasets are based on human-annotated noise, they model noisy real-world datasets more realistically, compared to synthetic data alternatives. The training set for both datasets contains 50,000 image-label pairs, and the test set contains 10,000 image-label pairs that are free from noise. For CIFAR-10N we utilize three noise settings for experiments in the paper– (1) Worst, which is the dataset version with the highest noise rate (40.21%) as the worst possible annotation label for the image is chosen, (2) Aggregate, which is the least noisy dataset (9.03%) as labels are chosen via majority voting amongst the annotations, and (3) Random which has intermediate noise (17.23%) and consists of picking one of the annotators’ labels. We use the first annotator for the random labels. For CIFAR-100N there is only a single noisy setting (Noisy100) due to the large number of labeling classes, and the overall noise rate is 40.20%.
B.1.3 GLUE Datasets
@@
The GLUE or the General Language Understanding Evaluation [Wang et al., 2018] benchmark datasets consist of a number of benchmarks for training, evaluating, and analyzing natural language models. As in the DataInf paper [Kwon et al., 2024], we utilize the four binary classification subset datasets: QNLI, SST2, QQP, and MRPC for experiments. Here, these datasets cover a wide variety of natural language task domains. For instance, QNLI [Wang et al., 2018] covers natural language inference, SST2 [Socher et al., 2013] covers sentiment analysis, QQP  [^4] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs covers question answering, and MRPC [Dolan and Brockett, 2005] covers paraphrase detection. We use the same datasets as in Kwon et al. [2024], where the training and test splits are obtained from the Huggingface datasets [^5] https://huggingface.co/docs/datasets library. For QQP and SST2 in Kwon et al. [2024] 4500 training samples and 500 test samples were randomly sampled from the full sets, so we utilize these in our experiments for a fair comparison.
B.1.4 Sentence Transformations
@@
For this benchmark dataset proposed in Kwon et al. [2024], the LLM is required to perform a specific transformation on an input sentence. There are 10 different sentence transformations. To help the model learn different transformations, “chatbot” name identifiers are used and each is uniquely associated with each transformation. These are the categories of sentence transformations (taking an example input sentence as “Welcome to the real world.”):
@@
B.1.5 Math With/Without Reasoning
@@ B.2 Models and Methods
We now describe the models and the methods used in our experiments throughout the main paper. First, we describe the ResNet-34 [He et al., 2016] architecture used as the base model for the noisy vision datasets, then the RoBERTa [Liu et al., 2019] NLP transformer model, and then the Llama-2 LLM. [^6] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf. We also describe implementation details and parameter values for the label correction baselines in Sections 4 and 5 and the influence-based baselines used throughout the paper. Finally, we also describe some key implementation details regarding our outlier gradient analysis approach.
B.2.1 ResNet-34
@@
The ResNet-34 model was proposed in He et al. [2016] and is a 34-layer convolutional neural network pretrained on the ImageNet-1K dataset at resolution 224 $\times$ 224. The pretrained model block is fine-tuned on the CIFAR-10N/CIFAR-100N training set experiments with default parameters– minibatch size (128), optimizer (SGD), initial learning rate (0.1), momentum (0.9), weight decay (0.0005), and number of epochs (100), for all experiments. Moreover, we directly used the implementation provided by Wei et al. [2022] and made modifications to their code.
B.2.2 RoBERTa
@@
As in Kwon et al. [2024], we utilize LoRA fine-tuning to fine-tune the RoBERTa-large model, a 355M parameter transformer language model that improves upon the original BERT model in key ways such as implementation and hyperparameter selection. LoRA is applied to every value matrix of the attention layers of the RoBERTa model. The pre-trained model from Huggingface is used. [^7] https://huggingface.co/docs/transformers/model_doc/roberta. A learning rate of 0.0003 and a batch size of 32 is used. The model is fine-tuned over 10 epochs using LoRA and dropout is set to be 0.05 while the rank of the LoRA matrix is set to 4, as recommended in Kwon et al. [2024]. The loss function used is a negative log-likelihood as the datasets are all for binary classification. The LoRA training is enabled using the Huggingface PEFT library. [^8] https://huggingface.co/docs/peft/index. For the influence experiments we have utilized the code provided in Kwon et al. [2024] and adapted it for our experiments. Moreover, we only compute influences using the training set gradients, and keep the test set hidden from the learning model for fair evaluation.
B.2.3 Llama2-13B-chat LLM
@@
We fine-tune the Llama2 13B parameter instruction tuned LLM using LoRA fine-tuning (applied to every query and value matrix of the attention layer) as in Kwon et al. [2024]. The LoRA parameters are as follows: learning rate is set to be 0.0003, rank of LoRA matrix is set to 8, $\alpha=32$ in 8-bit quantization, and the batch size is set to 32 across 25 fine-tuning epochs. A negative log-likelihood of the generated response is used as the loss function for fine-tuning as before. Here too, we adapt the code provided by Kwon et al. [2024] for our use cases.
B.2.4 Label Correction Baselines
@@
For label correction baselines in Sections 4 and 5Normalized Margin [Northcutt et al., 2021], Self-Confidence [Müller and Markert, 2019], and Confidence-Weighted Entropy [Kuan and Mueller, 2022], we utilize the implementation provided in the Cleanlab [^9] https://github.com/cleanlab/cleanlab/. library. We use default parameters for all three baselines. Note that the baselines are model agnostic and only require predicted labels and associated probabilities for predictions, which we can easily obtain from classifiers.
B.2.5 Influence-Based Baselines
@@
B.2.6 Outlier Gradient Analysis
We now discuss implementation details regarding outlier gradient analysis. Owing to the simplicity of our approach, the implementation is straightforward and follows directly from the algorithm. In most cases, we directly utilize the gradients obtained from the last layer of the model being considered. However, in some cases, the gradient space of samples can be high dimensional. For instance, for CIFAR-100N, the gradient space is of dimension 50000 $\times$ 51200 which unnecessarily increases memory and time complexity of outlier detection. As a result, we reduce the gradient space dimensionality by employing a sparse random projection step [Li et al., 2006] where the reduced dimension is ascertained using the scikit-learn library. We also utilize sparse random projection in this manner for the Llama-2-13B-chat LLM model experiments to reduce the dimensionality of the gradient space obtained.
Appendix C Additional Results and Experiments
We now provide details on additional experiments. We first provide results for the noisy label datasets and vision models shown in the main paper, but with standard deviation included. Then we conduct ablation experiments on the outlier detection threshold $k$ for the outlier gradient analysis algorithm. We also provide experiments on running time of our proposed approach (as well as details on computational complexity), ablation experiments on varying iForest parameters, results on ImageNet, experiments with ResNet-18 as the base model instead of ResNet-34, among others.
@@ Appendix C Additional Results and Experiments
Normalized Margin [Northcutt et al., 2021] 91.33 ± 0.11 90.06 ± 0.14 83.57 ± 0.32 60.94 ± 0.59
Self-Confidence [Müller and Markert, 2019] 91.38 ± 0.19 90.09 ± 0.17 83.65 ± 0.21 60.51 ± 0.51
Confidence Entropy [Kuan and Mueller, 2022] 91.11 ± 0.34 90.05 ± 0.26 83.63 ± 0.41 60.62 ± 0.26
Gradient Tracing [Pruthi et al., 2020] 91.47 ± 0.21 89.98 ± 0.20 83.38 ± 0.58 60.73 ± 0.38
LiSSA [Koh and Liang, 2017] 91.49 ± 0.34 90.05 ± 0.31 83.38 ± 0.58 60.48 ± 0.29
DataInf [Kwon et al., 2024] 91.46 ± 0.17 90.05 ± 0.38 83.40 ± 0.56 60.70 ± 0.31
Self-LiSSA [Bejan et al., 2023] 92.07 ± 0.15 89.58 ± 0.11 83.01 ± 0.34 59.48 ± 0.43
@@
C.1 Full Results with Standard Deviation for Vision Model Experiments
In the main paper results of Section 5 we provide accuracy values without the standard deviation listed, due to space constraints. Here, we augment those results by also providing the standard deviation obtained over the 5 runs. These results are denoted in Table 4. It can be seen that the standard deviations are in general low, and overall, outlier gradient trimming has low variance.
C.2 Additional Results for Different Trimming Budget $k$
We now conduct experiments varying $k$ from 2.5% to 12.5% for all three noise settings and baselines in the CIFAR-10N dataset. These results are shown in Table 5. As can be observed, our outlier analysis approaches features in the top-2 irrespective of the value of $k$. Moreover, the highest values across each noise regime are obtained by outlier gradient analysis (L2 norm at 12.5% for Aggregate and Random; and L2 norm at 2.5% for Worst). Finally, we find that setting $k$ as 5% and 12.5% are good overall choices leading to consistently desirable performance. Hence, we select 5% as the outlier budget in experiments.
Table 5: Varying the trimming budget $k$ and measuring test set performance across noisy datasets (top-2 performers at each $k$ in bold).
CIFAR10N (Aggregate) 2.5% 5% 7.5% 10% 12.5%
Gradient Tracing 92.11 91.47 92.17 91.99 91.98
LiSSA 92.08 91.49 91.83 92.27 91.74
DataInf 92.34 91.46 91.81 91.80 92.07
Self-LiSSA 91.71 92.07 91.32 91.72 91.33
Self-DataInf 91.22 91.41 91.37 91.29 91.15
Outlier Gradient (L1) 91.39 91.86 92.05 92.36 92.21
Outlier Gradient (L2) 92.10 92.21 92.70 92.63 92.78
Outlier Gradient (iForest) 91.77 91.36 91.57 91.92 92.08
CIFAR10N (Random) 2.5% 5% 7.5% 10% 12.5%
Gradient Tracing 90.71 89.98 90.41 90.75 90.96
LiSSA 90.21 90.05 91.09 90.88 90.00
DataInf 90.77 90.05 90.30 90.26 90.80
Self-LiSSA 89.76 89.58 89.50 88.94 89.49
Self-DataInf 89.91 89.81 90.32 89.91 90.00
Outlier Gradient (L1) 90.51 90.66 90.24 90.45 91.17
Outlier Gradient (L2) 90.72 90.25 90.63 90.50 91.21
Outlier Gradient (iForest) 90.03 90.20 90.06 90.38 90.62
CIFAR10N (Worst) 2.5% 5% 7.5% 10% 12.5%
Gradient Tracing 83.56 83.38 83.61 84.12 84.49
LiSSA 84.51 83.38 84.25 83.63 83.89
DataInf 84.31 83.40 83.45 84.01 84.12
Self-LiSSA 82.65 83.01 82.75 82.71 82.66
Self-DataInf 83.70 83.15 83.53 82.96 83.84
Outlier Gradient (L1) 84.26 84.20 84.12 84.32 84.25
Outlier Gradient (L2) 84.48 82.99 84.09 84.35 84.43
Outlier Gradient (iForest) 83.74 83.72 84.22 84.44 83.25
C.3 Experiments on Running Time and Computational Complexity
We now present running time experiments for outlier gradient analysis on both the CIFAR-10N and CIFAR-100N datasets compared to the other baselines compared in the paper in Table 6. It can be seen that outlier gradient analysis is computationally efficient and a fraction of the original running time of the model. Moreover, it is order of magnitudes faster than the other baselines. Thus, our outlier gradient analysis approach is computationally efficient as an option for trimming detrimental samples and improving model performance. Most notably, only Gradient Tracing is faster than outlier gradient analysis, but as we demonstrated in the main paper results, it seldom as accurate in detecting detrimental samples as outlier analysis. Thus, outlier gradient analysis is ideal for balancing performance with computational efficiency. We also provide analytical time complexity comparisons in Table 7. Although, it is important to note that in practice, outlier gradient analysis is much faster than the worst case time complexity, as can be seen in Table 6.
Table 6: Running time for our outlier gradient analysis approaches and other baselines (top-2 in bold).